A father is twice as old as his son
2.1if the sons age is represented by x how old is the father
2.2how old were the father and son 20years back
2.3 20years ago the father was six times his sons age represent this as an equation

Answer by addingup(3677)

So the father is twice, meaning two times, older than the son. In algebraic terms: f = 2s Since the problem says that x stands for the son's age, substitute s with x and you have your answer.
f-20 = 2s-20
f-20 = 6(son-20)
Remember to substitute for x